OpenStudy (anonymous):

Suppose you are playing a game with two number cubes. Let A represent rolling 2, 3, or 4, and B represent rolling 1, 5, or 6. The probability of A is 1/2 and the probability of B is 1/2 . a. Simplify (1/2 a + 1/2 b) ^2 b. What is the probability that one number cube shows 2, 3, or 4, and the other shows 1, 5, or 6?

OpenStudy (anonymous):

1. Use: \[(a + b)^2 = a^2 + b^2 + 2ab\] \[= \frac{a^2}{4} + \frac{b^2}{4} + \frac{a \times b}{2}\] As a = 1/2 and b = 1/2 \[= \frac{1}{16} + \frac{1}{16} + \frac{1}{8} = \frac{1 + 1 + 2}{16} = \frac{4}{16} = \frac{1}{4}\]

OpenStudy (anonymous):

b> \[P = \frac{1}{3} \times \frac{1}{3} = \frac{1}{9}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome dear..